Accelerator Update for August 1, 2008 - August 4, 2008
Friday August 1
The day shift began with Operations monitoring the Tevatron (TeV) store 6326 (104.7E30), the Recycler (RR) stash (179.2E10) and the Antiproton Source stack (30.3E10), with Linac off for RF station (LRF4) repairs, with the Main Injector and Booster on with no beam, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, and with no beam to the SciBooNE / MiniBooNE (BNB), NuMI, and Meson MTest experiments.
The Cryo Coordinator reported at 9:39 AM that an expert had reset the thermometry crate at TeV sector C2.
Operations reported at 12:54 PM that Linac experts had completed their LRF4 repairs.
Linac experts reported at 1:27 PM that they were shutting off a Klystron RF station (KRF3) to make sure its regulation flow valve was operating properly.
Linac experts brought LRF3 on line at 2:22 PM.
Pbar resumed stacking and BNB resumed taking beam at 2:29 PM
Operations disabled the MTest beam at 2:37 PM
At 2:38 PM, Operations reported that the NuMI beam permit had been tripping due to RF problems. An RF expert investigated.
Operations terminated store 6326 at 3:16 PM.
Operations began preparing the D0 collision hall for access at 3:27 PM.
Operations reported at 4:02 PM that the Recycler had stashed antiprotons four times during the day shift.
The Cryo Coordinator reported a problem with the TeV C4 cryo system at 4:06 PM. Experts investigated.
D0 came out of access at 5:26 PM.
At 5:57 PM, MIRF14 tripped off and took down the NuMI beam permit. Everything reset.
Operations established beam to MTest at 8:23 PM.
At 10:10 PM, Operations reported that a TeV D0 quadrupole (D0Q6) power supply wasn’t ramping. Operators changed a control card and performed a TECAR restore of all the D0 low beta magnet control cards, but this did not help. Operations called in a power supply expert.
Operations reported at 11:58 PM that the Recycler had stashed four times during the evening shift.
Saturday August 2
The midnight shift began with TeV power supply experts working on D0Q6, with Operations monitoring the stash (392.9E10), with Pbar stacking (72.5E10), with the BNB, NuMI, and MTest experiments taking beam.
At 3:52 AM, Operations reported that Pbar was losing antiprotons. This lasted about five minutes. 7E10 were lost.
At 5:23 AM, LRF4 tripped off for the first of two times in four minutes. It reset.
TeV power supply experts reported at 6:29 AM that D0Q6 was back on line. A bad reference cable was found and reterminated.
Operations put the TeV through a dry squeeze at 6:49 AM.
Operations put the TeV through a wet squeeze at 7:22 AM.
Operations put the TeV through another wet squeeze at 8:40 AM.
Operations put the TeV into shot setup at 9:36 AM.
At 9:50 AM, the Linac low energy Buncher tripped off. It reset.
At 10:56 AM, the TeV suffered a sector F3 quench during the ramp. Operators and experts investigated.
Operations established beam to MTest at 11:57 AM.
At 2:29 PM, LRF4 tripped off and would not reset. Operations called in an expert.
The MTest experiment reported at 2:40 PM that they were done with this part of their experiment. Operations disabled the MTest beam line. A new MTest user will begin on August 13, 2008.
At 3:04 PM, the TeV sector C3 electronics room air conditioner tripped off. The Duty Mechanic investigated and installed a portable AC unit.
At 3:05 PM, Operations reported more problems at the TeV sector C3 electronics room; the VFCs were not reading back. Operations changed out the VFC chassis.
At 3:26 PM, the Linac Klystron Buncher tripped on a LCW flow indication.
At 4:42 PM, Operations reported that the heater at TeV sector F4 did not fire properly. Operators changed the HFU, but that didn’t help. HFU-4A was still bad. Operations contacted experts.
A Linac expert reported at 6:44 PM that LRF4 was operating normally. Operations established beam to Pbar, NuMI, and BNB.
Operations reported at 8:01 PM that the Recycler had stashed three times during the last 12 hours.
TeV experts reported at 8:44 PM that the TeV sector F4 HFU4 A2 problem was due to an open contact in the tunnel. Operators prepared the TeV for an access.
At 9:48 PM, BRF8 tripped off. It reset.
Around 10:18 PM, Operations reported a Recycler vacuum burst. About 20E10 of antiprotons were initially lost. Operations contacted experts.
TeV experts reported at 10:41 PM that the HFU cables were fine and, unfortunately, the problem went away. Operators conducted a TeV hipot after the TeV experts came out of access; everything looked good.
Per Recycler expert instruction at 10:58 PM, operators turned on vacuum pumps and opened beam valves. About 150E10 of antiprotons in total were lost.
Operations reported at 11:01 PM that cryo system experts were checking out a heat load problem for the TeV sector D1 cryo system.
Sunday August 3
The midnight shift began with the TeV recovering from the quench, with Operations monitoring the stash, with Pbar stacking, and with the BNB and NuMI experiments taking beam.
At 12:48 AM, CDF reported that they had a read out crate fan failure.
Operations prepared CDF for access at 1:26 AM. They came out of access about 45-minutes later.
At 1:37 AM, the TeV sector C3 electronics room temperature went high. Operators investigated and found that the portable AC unit’s evaporator had frozen. Operators put the unit into fan only mode until the ice ball melted.
Operations reported at 2:49 AM that the Recycler emittances had recovered from the vacuum burst.
At 2:50 AM, Operations reported that the TeV sector D1 cryo problems continued. Cryo experts reported that they might have to replace the cold compressor.
Cryo system experts at CHL reported at 3:56 AM that they were turning off the cold compressor at D1 to begin its replacement.
At 5:24 AM, KRF4 went out of phase. Operators rebooted the Linac low level RF system.
Cryo system experts reported at 6:32 AM that the D1 cryo cooldown was beginning.
Operations reported at 8:02 AM that the Recycler had stashed seven times during the last 12 hours.
At 9:10 AM, Operations reported that the NuMI beam permit tripped off for the first of two times in about 45 minutes on a RF problem, but no stations were tripped. Later, operators discovered that MIRF17 had a weak Power Amplifier.
At 10:03 AM, LRF4 tripped off. It reset.
At 10:05 AM, Operations reported that the C3 portable AC unit froze again. The Duty Mechanic investigated and got it going.
Cryo experts reported at 10:46 AM that the sector D1 cold compressor pump down had began. It finished an hour later.
CHL reported at noon that they were having a problem with one of their big compressors (Comp D). They brought up a backup compressor.
At 12:15 PM, Operations reported that a problem with the TeV sector F4 QPM was stopping them from turning on the TeV. Operators investigated and found inductance problems.
At 12:55 PM, Operations reported that the I- Source was tripping off frequently, but it would come right back on.
At 2:41 PM, Operations reported that LRF2 was causing beam position problems.
Operations reported at 4:26 PM that they had completed several QBS ramps with no problems. Experts took the TeV sector F2 power supply out of the circuit.
At 5:27 PM, the 400MeV area beam permit tripped off four times in three minutes due to LRF2 problems. It reset.
Operations put the TeV through a dry squeeze at 5:31 PM.
Operations put the TeV into shot setup at 7:41 PM.
Operations reported at 8:03 PM that the Recycler had stashed two times during the last 12 hours.
Operations established store 6334 at 9:13 PM with an initial luminosity of 273.9E30.
At 10:02 PM, Operations reported more NuMI trips due to MIRF17. Operations lowered NuMI’s beam intensity.
Monday August 4
The midnight shift began with Operations monitoring store 6334 and the stash, with Pbar stacking, and with the BNB and NuMI experiments taking beam.
At 6:43 AM, the Pelletron tripped off and would not reset. Operations contacted an expert.
At 7:42 AM, the TeV suffered a sector C3 quench due to a lightening strike on the C3 service building. Store 6334 was lost. A witness from D0 actually saw the building struck.
Operations reported at 8:04 AM that the Recycler had stashed seven times during the last 12 hours.
The Plans for Monday
The plans for today are to repair any problems caused by the lightning strike of the C3 service building and then establish a new store.
Comments and Suggestions
|last modified 08/04/2008 email Fermilab|
|Security, Privacy, Legal|